What is: 2 1/6 divided by 3 1/3 times 6 1/4 divided by 4 1/3?
Answer by CharlesG2(828)

What is: 2 1/6 divided by 3 1/3 times 6 1/4 divided by 4 1/3?

2 1/6 = 12/6 + 1/6 = 13/6 (12/6 = 2)
3 1/3 = 9/3 + 1/3 = 10/3 (9/3 = 3)
6 1/4 = 24/4 + 1/4 = 25/4 (24/4 = 6)
4 1/3 = 12/3 + 1/3 = 13/3 (12/3 = 4)

13/6 divided by 10/3 = 13/6 * 3/10
(when you divided 2 fractions you multiply by the reciprocal of the 2nd fraction)
13/6 * 3/10 = 39/60 = 13/20

13/20 times 25/4 = 13/20 * 25/4 = 325/80 = 65/16
(multiplied straight across then simplified)

65/16 divided by 13/3 = 65/16 * 3/13
65/16 * 3/13 = 195/208
195/208 = 0.9375 in decimal form --> this is your answer